OpenStudy (anonymous):

Given two rectangles, one with length = l and width = w, and the second rectangle with length = 2l and width = 4w. How much larger will the perimeter of the second rectangle be than the perimeter of the first rectangle?

OpenStudy (anonymous):

Perimeter1 = 2(l + w) Perimeter2 = 2(2l + 4w) Subtract P1 from P2